Find the slope of any line perpendicular to the line through points (0, 5) and (-3,-4).
Answer by stanbon(75887)

Find the slope of any line perpendicular to the line through points (0, 5) and (-3,-4).
---------
slope thru the given points: [5--4]/[0--3] = 9/3 = 3
-------------
Any line perpendicular to the given line must have slope = -1/3
Cheers,
Stan H.